OpenStudy (anonymous):

simplify the expression 2^5*3^4/2^7*3^2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\huge \frac{2^53^4}{2^73^2}\] \[\huge =\frac{3^{4-2}}{2^{7-5}}\] is a good start

OpenStudy (anonymous):

DON'T YOU HAVE TO GET RID OF THE NEGATIVES